# TNO, The Hague, The Netherlands

### v1

@@ 3.6 Prediction and Inventory Control
The next category of application is that of demand prediction. There are many papers on Quantum Machine Learning (QML) and prediction, however only one on a direct application within logistics and supply chain optimisation. Jahin et al. jahin2023qamplifynet focus on improving backorder prediction in supply chain management (SCM) through the introduction of a novel hybrid quantum-classical neural network called QAmplifyNet. Backorder prediction is used for optimising inventory control, reducing costs, and enhancing customer satisfaction. Traditional machine-learning models face challenges with large-scale datasets and complex relationships. The proposed QAmplifyNet demonstrates high accuracy in predicting backorders, particularly on short and imbalanced datasets, outperforming classical models and quantum neural networks. The study also addresses the scarcity of research on product backordering, emphasising the challenges of class imbalance in datasets. While classical machine learning models have been widely used, following the authors, quantum-inspired techniques show superior performance, specifically QAmplifyNet. The research integrates quantum and classical computing to leverage the strengths of both, overcoming challenges in handling short, imbalanced datasets common in SCM. The methodological framework involves evaluating seven preprocessing techniques, selecting the best-performing one, and enhancing the model’s interpretability using Explainable Artificial Intelligence (XAI) techniques. Practical implications include improved inventory control, reduced backorders, and enhanced operational efficiency. QAmplifyNet achieves a high F1-score of 94% for predicting “Not Backorder” and 75% for predicting “Backorder”, along with the highest AUC-ROC score of 79.85%. The hybrid quantum-classical model shows promise for real-world SCM applications, paving the way for further exploration of quantum-inspired machine learning in supply chain management.
@@
The work by Jiang jiang2022quantum highlights the challenges in classic computing related to large state and action spaces in supply chain problems. They introduce a quantised policy iteration algorithm designed for inventory control problems and demonstrate its effectiveness. They perform simulations and experiments using IBM Qiskit and the qBraid system. They underscore the practicality of variational algorithms for solving small-sized inventory control problems.
@@ 3.7 Scheduling
In this section, we delve into the application of quantum algorithms to address scheduling problems within the domains of logistics and supply chain optimisation. Scheduling plays an important role in orchestrating the efficient flow of resources, managing timelines, and optimising overall operational processes. The research by Riandari et al. riandari2021quantum explores the potential of quantum computing in production planning or scheduling, using quantum algorithms like QAOA and VQE to minimise production costs and address demand and resource constraints. They evaluate the potential benefits, limitations, and difficulties associated with employing quantum computing to production planning.
@@
The work by Ajagekar et al. ajagekar2020quantum ; ajagekar2020quantum2 discusses multiple optimisation problems, including the job-shop scheduling problem (JSP). They propose a hybrid approach to address this complex problem, which involves determining an optimal schedule for a set of jobs on a set of machines. The proposed MILP model incorporates due dates and sequence-independent processing times, with decision variables representing start times, assignments, and sequencing of jobs. The hybrid QC-MILP decomposition method involves two phases: solving the MILP problem with a classical Gurobi solver and employing QC for sequencing. The quantum processor, QA, is utilised to determine feasible schedules, overcoming memory limitations of classical solvers for larger instances. Computational results indicate competitive performance, particularly for large-scale industrial problems, where classical solvers face limitations. The approach says to guarantee a global optimal solution but may require longer runtimes, showcasing the complementary strengths of MILP and QC in tackling intricate scheduling problems effectively. The other problem they study is the Manufacturing Cell Formation (MCFP) problem. Cellular manufacturing involves organising equipment to facilitate continuous flow production, enhancing efficiency. Here, they propose a hybrid QC and Mixed-Integer Quadratic Programming (MIQP) approach to solve the MCFP. The MIQP model is formulated to minimise costs associated with intracellular movement, resource utilisation, and machine set-ups. The dual LP model is solved classically, and its results guide the quantum step, solving a QUBO problem. The hybrid method efficiently handles large instances, outperforming conventional solvers. The proposed approach provides optimal or near-optimal solutions, demonstrating again the complementary strengths of classical and quantum computing in tackling complex manufacturing problems.
@@
The article by Permin et al. permin2022simple explores the application of quantum computers to simplify the complexity of Job Shop Scheduling (JSP). The study employs Grover’s algorithm in IBM Qiskit to translate a JSP into a binary qubit format, effectively reducing the search space. By conducting simulations, the research demonstrates a significant decrease in complexity
@@
The article by Venturelli et al. venturelli2015quantum also focuses on the JSP. Here, QA is employed as a solver for JSP, with a detailed presentation of the formulation and strategies used. The study discusses the challenges associated with quantum annealing, including limitations in precision, connectivity, and the number of variables. The quantum annealer’s performance is compared with classical global-optimum solvers, particularly, as the authors call them, algorithms B (based on Brucker et al. brucker1994branch ) and MS (based on Martin et al. martin1996new ). Results indicate that D-Wave 2 QA faces challenges in outperforming classical algorithms on small instances of JSP, raising questions about its scalability and asymptotic advantage. Motivated by these quantum annealing results, the study proposes a QA JSP solver, addressing issues such as embedding, pre-processing, and running strategies. The article introduces a solving process for a full JSP optimisation solver, inclusive adapting classical algorithms as pre-processing techniques. The proposed timespan discrimination is highlighted as a flexible compromise between full optimisation and decision formulations, allowing for immediate benefits from precision improvements.
@@
Also Kurowski et al. kurowski2020hybrid employ a QA approach for the JSP. The mapping of the JSP to the QUBO formulation is detailed, considering the constraints and the specific topology of the D-Wave 2000Q. The proposed heuristic is explained, involving the decomposition of the JSP into smaller optimisation problems and the tuning of QUBO parameters. Experimental studies on a JSP test instance (ft06) are presented, including estimates of the number of feasible solutions. The results obtained from the quantum annealing heuristic are compared with optimal solutions generated by classical heuristic methods for the JSP benchmark.
@@
The study by by Amaro amaro2022case explores the application of VQAs on IBM’s superconducting quantum processors to solve the JSP. The four VQAs tested are the QAOA, general VQE, variational quantum imaginary time evolution (VarQITE), and the recently introduced filtering VQE(F-VQE). The study compares convergence speed, scalability, and accuracy of the VQAs, with a focus on practical applicability to non-fault-tolerant hardware. The results show that F-VQE outperformed other algorithms in terms of convergence speed and frequency of sampling the optimal solution, even on instances with five qubits. VarQITE exhibited slower convergence but sampled optimal solutions more frequently. VQE converged slowly and sampled optimal solutions less frequently, while QAOA struggled with convergence due to the complexity of circuits and optimiser choice. In a subsequent set of experiments, F-VQE was further tested on larger instances with up to 23 qubits, showcasing its promising performance and ability to handle larger combinatorial optimisation problems. The study suggests that F-VQE is a step towards addressing challenges associated with quantum hardware limitations, including sparse connectivity and cross-talk noise.
@@
Schworm et al. schworm2023solving explore the application of QA as a metaheuristic for solving the JSP, particularly the flexible job shop scheduling problems (FJSP). QA utilises quantum mechanical effects to evaluate multiple solutions simultaneously, potentially providing time-efficient solutions for complex assignment problems. The study assesses QA’s efficiency, scalability, solution quality, and computing time by solving FJSP instances of various sizes. The results demonstrate QA’s potential in finding high-quality solutions for the FJSP within seconds, even for large instances, making it a promising approach for industry-scale applications. The evaluation includes a scientific benchmark comparing QA with state-of-the-art algorithms, highlighting QA’s efficiency in finding high-quality solutions in a short time. While QA shows promise, challenges such as dynamic factors in industrial settings and the need for multi-objective optimisation are acknowledged for future research. The study concludes that QA can be a valuable tool for intelligent transportation systems, emphasising the need for further improvement and exploration of QA’s capabilities.
@@
The paper by Denkena et al. denkena2021quantum introduces a new method for process-parallel Flexible Job Shop Scheduling (FJSP) using QA. The FJSP requires continuous and process-parallel optimisation of machine allocation and processing sequences. The proposed quantum annealing-based method demonstrates results comparable to classical heuristics with reliably low anneal times, especially for large problem instances. The approach is shown to be robust against stochastic influences and hyper-parameter variations. In a practical use case, the approach outperforms classical MES solutions and combinations of MES with genetic algorithms in terms of makespan reduction.
@@
This paper by Carugno et al. carugno2022evaluating also investigates the application of Quantum Annealing (QA) to the JSP. The study provides a detailed analysis of applying QA to JSP, covering problem formulation, quantum annealer fine-tuning, and comparison with classical solvers. The research addresses often-overlooked aspects, such as computational costs, qubit requirements, and mitigating chain breaks. Additionally, advanced tools like reverse annealing are explored. Results reveal challenges at various stages, highlighting important research questions and areas for improvement in quantum annealing technology. The study contributes to a balanced discussion on QA’s potential and limitations in solving JSP, emphasising open research directions for further exploration.
@@
A factory scheduling problem was presented by Yarkoni et al. yarkoni2021multi , derived from a real-world optimisation problem involving re-ordering colour sequences in a paint shop during the manufacturing of cars. The task is to assign colours to sequences of cars such that all customer orders are fulfilled while simultaneously minimising the number of colour switches within the sequence. In this work, the authors investigated problem instances ranging from 10–3000 cars in a single sequence. The multi-car paint shop problem described in the paper has a simple Ising model representation, and instances derived directly from real-world data were solved using D-Wave’s QPUs, hybrid algorithms, and classical algorithms. The results showed that while the quantum hardware and hybrid algorithms were able to provide adequate solutions to the problem for smaller sizes, the performance of these algorithms approached that of a simple greedy algorithm in the large size limit. However, the analysis also concluded that QA is currently approaching the limit of industrially-relevant problem scales. Partly the same authors look at the same problem but then using the QAOA approach in streif2021beating . This research combines numerical simulations with experimental data obtained from a trapped ion quantum computer, presenting a comparison between the performance of QAOA and classical heuristics, particularly in the infinite-size limit for noiseless quantum computation. The experimental findings show the quick degradation of the quantum algorithm’s efficiency as the problem size increases.
@@
Also Huang et al. huang2022paint look at the paint shop problem. While previous research primarily targeted minimising color changeovers between vehicles, this study introduces a novel problem formulation considering not only color transitions but also the impact on product quality. By incorporating a machine learning model, which is not explained in the paper, to predict quality defects and leveraging quantum computing, the paper presents an approach for solving the generalised paint shop sequencing problem. The paper concludes with a case study showcasing the potential of QAOA in solving a small-scale sequencing problem, paving the way for future exploration in real-world, industrial-sized scenarios.
@@
The next problem coming from the aircraft industry, is the Aircraft Recovery Problem (ARP), a crucial aspect of airline disruption management. Given the financial pressures on airlines, aggravated by the recent pandemic, a swift and efficient solution to recover from disruptions is essential to minimise financial losses. Mori et al. mori2022replanning develop a Two-Stage algorithm to model ARP as a QUBO problem. Two experiments are conducted to evaluate the model’s performance using QC in comparison to classical solvers. The first experiment compares the proposed approach with a previous one de2015evolutionary based on evolutionary computation algorithms, showing similar execution times and results. The second experiment compares a classical solver with a hybrid solver using the proposed model, demonstrating identical solutions but with faster execution times for the hybrid solver.
@@
A more generic problem that also has applications in the aircraft industry is the Quadratic Assignment Problem (QAP). Mohammadbagherpoor mohammadbagherpoor2021exploring explores the QAP in the application of airline gate-scheduling . Here, the QAP involves optimising the allocation of facilities (airline gates) to locations (flights) with known assignment costs. The paper introduces the concept of using VQE and space-efficient graph colouring, to enhance quantum computing algorithms for solving QAPs. The study tests these enhanced quantum computing algorithms on an 8-airline gate and 24-flight test case using both the IBM quantum computing simulator and a 27-qubit superconducting transmon IBM quantum computing hardware platform. The results suggest that smaller circuit size and efficient mapping to hardware topology enable the application of variational quantum algorithms to larger problem sets, a critical advancement in solving optimisation problems.
@@
The work by Scherer et al. scherer2021oncall discusses the application of Grover’s algorithm, to solve the “on-call spacecraft operator scheduling” problem at the German Space Operating Center (GSOC). The problem involves scheduling operators for spacecraft missions, considering various constraints and optimisation goals. The scheduling must allocate approximately 50 operators to 20 positions over 180 days, while satisfying constraints such as operator availability, working preferences, and limiting consecutive working days. Evaluation is done using Qiskit’s QASM simulator with up to 32 qubits. The success rates are evaluated for different problem sizes and configurations. The paper successfully demonstrates solving the scheduling problem using Grover’s algorithm on a quantum simulator. Future work includes testing on real quantum devices, improving qubit stability, and exploring quantum-classical hybrid approaches. Emphasis on minimising the number of qubits and optimising the algorithm’s width and depth is highlighted.
@@
Adelomou et al. adelomou2020using address personnel scheduling. They look at the challenges faced by social workers in generating optimal visiting schedules for patients. The problem involves a combination of routing and planning issues. The article suggest the use of the Parameterised Quantum Circuits (PQC) and the VQE algorithm in solving the problem. It explores the potential of PQC as a basis for Machine Learning (ML) and delves into the formulation of the social worker problem for quantum computing. The proposed methodology aims to offer an efficient and intelligent system, capable of learning by analogy and avoiding the need for recalculating from scratch. Experimental results suggest the effectiveness of the Quantum Cases-Based Reasoning (QCBR) approach, demonstrating potential benefits for optimising worker schedules. The article concludes by emphasising the broader applicability of the proposed formulation to solve various planning, scheduling, and routing problems. It envisions the creation of hybrid and intelligent systems that leverage quantum resources efficiently, reducing costs and laying the foundation for generalised problem-solving in the field.

### v2

@@ 3.6 Prediction and Inventory Control
The next category of application is that of demand prediction. There are many papers on Quantum Machine Learning (QML) and prediction, however only one on a direct application within logistics and supply chain optimisation. Jahin et al. [69] focus on improving backorder prediction in supply chain management (SCM) through the introduction of a novel hybrid quantum-classical neural network called QAmplifyNet. Backorder prediction is used for optimising inventory control, reducing costs, and enhancing customer satisfaction. Traditional machine-learning models face challenges with large-scale datasets and complex relationships. The proposed QAmplifyNet demonstrates high accuracy in predicting backorders, particularly on short and imbalanced datasets, outperforming classical models and quantum neural networks. The study also addresses the scarcity of research on product backordering, emphasising the challenges of class imbalance in datasets. While classical machine learning models have been widely used, following the authors, quantum-inspired techniques show superior performance, specifically QAmplifyNet. The research integrates quantum and classical computing to leverage the strengths of both, overcoming challenges in handling short, imbalanced datasets common in SCM. The methodological framework involves evaluating seven preprocessing techniques, selecting the best-performing one, and enhancing the model’s interpretability using Explainable Artificial Intelligence (XAI) techniques. Practical implications include improved inventory control, reduced backorders, and enhanced operational efficiency. QAmplifyNet achieves a high F1-score of 94% for predicting “Not Backorder” and 75% for predicting “Backorder”, along with the highest AUC-ROC score of 79.85%. The hybrid quantum-classical model shows promise for real-world SCM applications, paving the way for further exploration of quantum-inspired machine learning in supply chain management.
@@
The work by Jiang et al. [71] highlights the challenges in classic computing related to large state and action spaces in supply chain problems. They introduce a quantised policy iteration algorithm designed for inventory control problems and demonstrate its effectiveness. They perform simulations and experiments using IBM Qiskit and the qBraid system. They underscore the practicality of variational algorithms for solving small-sized inventory control problems.
A recent book provides some new insides. The chapter by Sehrawat [140] explores the application of quantum machine learning (QML) for demand prediction in supply chain networks. Traditional demand forecasting methods often struggle with the complexities and uncertainties of modern supply chains. QML, leveraging quantum computing’s computational power and the adaptability of machine learning algorithms, offers significant advantages over classical methods. By applying QML techniques, organizations can improve demand prediction accuracy, optimize inventory, and enhance overall supply chain performance. Through case studies and practical examples, the chapter demonstrates how QML can support informed decision-making and drive improvements in supply chain management. The chapter by Gutta et al. [60] explores the integration of artificial intelligence (AI) and quantum machine learning (QML) to enhance supply chain forecasting. The combination of AI’s ability to analyse large datasets and extract insights with QML’s capacity for processing complex probabilistic distributions offers unparalleled accuracy in demand forecasting, inventory optimization, and risk mitigation. By leveraging these technologies, organizations can transform their forecasting processes, improve operational efficiency, and enhance competitiveness. The chapter discusses the synergistic potential of AI-infused QML models, demonstrating their impact through case studies and real-world applications in supply chain management. Last, the chapter by Koushik et al. [78] investigates the application of deep learning strategies to advance predictive maintenance within supply chain systems. Predictive maintenance plays a critical role in ensuring the reliability and efficiency of equipment and machinery across supply chain operations. Traditional approaches often struggle to effectively capture complex patterns and anticipate impending failures. However, deep learning techniques offer a promising solution by leveraging neural networks to analyze vast amounts of sensor data and identify early indicators of potential equipment malfunctions. This chapter explores various deep learning architectures, such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs), tailored for predictive maintenance applications. Through case studies and real-world examples, the chapter illustrates how deep learning strategies can empower organizations to proactively manage asset health, minimize downtime, and optimize maintenance schedules, thereby enhancing overall supply chain resilience and performance.
@@ 3.7 Scheduling
In this section, we delve into the application of quantum algorithms to address scheduling problems within the domains of logistics and supply chain optimisation. Scheduling plays an important role in orchestrating the efficient flow of resources, managing timelines, and optimising overall operational processes. The research by Riandari et al. [122] explores the potential of quantum computing in production planning or scheduling, using quantum algorithms like QAOA and VQE to minimise production costs and address demand and resource constraints. They evaluate the potential benefits, limitations, and difficulties associated with employing quantum computing to production planning.
@@
The work by Ajagekar et al. [3, 4] discusses multiple optimisation problems, including the job-shop scheduling problem (JSP). They propose a hybrid approach to address this complex problem, which involves determining an optimal schedule for a set of jobs on a set of machines. The proposed MILP model incorporates due dates and sequence-independent processing times, with decision variables representing start times, assignments, and sequencing of jobs. The hybrid QC-MILP decomposition method involves two phases: solving the MILP problem with a classical Gurobi solver and employing QC for sequencing. The quantum processor, QA, is utilised to determine feasible schedules, overcoming memory limitations of classical solvers for larger instances. Computational results indicate competitive performance, particularly for large-scale industrial problems, where classical solvers face limitations. The approach says to guarantee a global optimal solution but may require longer runtimes, showcasing the complementary strengths of MILP and QC in tackling intricate scheduling problems effectively. The other problem they study is the Manufacturing Cell Formation (MCFP) problem. Cellular manufacturing involves organising equipment to facilitate continuous flow production, enhancing efficiency. Here, they propose a hybrid QC and Mixed-Integer Quadratic Programming (MIQP) approach to solve the MCFP. The MIQP model is formulated to minimise costs associated with intracellular movement, resource utilisation, and machine set-ups. The dual LP model is solved classically, and its results guide the quantum step, solving a QUBO problem. The hybrid method efficiently handles large instances, outperforming conventional solvers. The proposed approach provides optimal or near-optimal solutions, demonstrating again the complementary strengths of classical and quantum computing in tackling complex manufacturing problems.
@@
The article by Permin et al. [111] explores the application of quantum computers to simplify the complexity of Job Shop Scheduling (JSP). The study employs Grover’s algorithm in IBM Qiskit to translate a JSP into a binary qubit format, effectively reducing the search space. By conducting simulations, the research demonstrates a significant decrease in complexity
@@
The article by Venturelli et al. [155] also focuses on the JSP. Here, QA is employed as a solver for JSP, with a detailed presentation of the formulation and strategies used. The study discusses the challenges associated with QA, including limitations in precision, connectivity, and the number of variables. The quantum annealer’s performance is compared with classical global-optimum solvers, particularly, as the authors call them, algorithms B (based on Brucker et al. [24]) and MS (based on Martin et al. [93]). Results indicate that D-Wave 2 annealer faces challenges in outperforming classical algorithms on small instances of JSP, raising questions about its scalability and asymptotic advantage. Motivated by these QA results, the study proposes a QA JSP solver, addressing issues such as embedding, pre-processing, and running strategies. The article introduces a solving process for a full JSP optimisation solver, inclusive adapting classical algorithms as pre-processing techniques. The proposed timespan discrimination is highlighted as a flexible compromise between full optimisation and decision formulations, allowing for immediate benefits from precision improvements.
@@
Also Kurowski et al. [80] employ a QA approach for the JSP. The mapping of the JSP to the QUBO formulation is detailed, considering the constraints and the specific topology of the D-Wave 2000Q. The proposed heuristic is explained, involving the decomposition of the JSP into smaller optimisation problems and the tuning of QUBO parameters. Experimental studies on a JSP test instance (ft06) are presented, including estimates of the number of feasible solutions. The results obtained from the QA heuristic are compared with optimal solutions generated by classical heuristic methods for the JSP benchmark.
@@
The study by by Amaro [7] explores the application of VQAs on IBM’s superconducting quantum processors to solve the JSP. The four VQAs tested are the QAOA, general VQE, variational quantum imaginary time evolution (VarQITE), and the recently introduced filtering VQE(F-VQE). The study compares convergence speed, scalability, and accuracy of the VQAs, with a focus on practical applicability to non-fault-tolerant hardware. The results show that F-VQE outperformed other algorithms in terms of convergence speed and frequency of sampling the optimal solution, even on instances with five qubits. VarQITE exhibited slower convergence but sampled optimal solutions more frequently. VQE converged slowly and sampled optimal solutions less frequently, while QAOA struggled with convergence due to the complexity of circuits and optimiser choice. In a subsequent set of experiments, F-VQE was further tested on larger instances with up to 23 qubits, showcasing its promising performance and ability to handle larger combinatorial optimisation problems. The study suggests that F-VQE is a step towards addressing challenges associated with quantum hardware limitations, including sparse connectivity and cross-talk noise.
@@
Schworm et al. [137, 138] explore the application of QA as a metaheuristic for solving the JSP, particularly the flexible job shop scheduling problems (FJSP). QA utilises quantum mechanical effects to evaluate multiple solutions simultaneously, potentially providing time-efficient solutions for complex assignment problems. The study assesses QA’s efficiency, scalability, solution quality, and computing time by solving FJSP instances of various sizes. The results demonstrate QA’s potential in finding high-quality solutions for the FJSP within seconds, even for large instances, making it a promising approach for industry-scale applications. The evaluation includes a scientific benchmark comparing QA with state-of-the-art algorithms, highlighting QA’s efficiency in finding high-quality solutions in a short time. While QA shows promise, challenges such as dynamic factors in industrial settings and the need for multi-objective optimisation are acknowledged for future research. The study concludes that QA can be a valuable tool for intelligent transportation systems, emphasising the need for further improvement and exploration of QA’s capabilities.
The same authors [139] present a Quantum Annealing-based Solving Algorithm for multi-objective flexible job shop scheduling. It optimizes makespan, total workload, and job priority simultaneously. The method decomposes large problems into subproblems, uses bottleneck factors for job selection, and demonstrates superior performance compared to classical algorithms in terms of solution quality and computation time.
@@
The paper by Denkena et al. [42] introduces a new method for process-parallel Flexible Job Shop Scheduling (FJSP) using QA. The FJSP requires continuous and process-parallel optimisation of machine allocation and processing sequences. The proposed QA-based method demonstrates results comparable to classical heuristics with reliably low anneal times, especially for large problem instances. The approach is shown to be robust against stochastic influences and hyper-parameter variations. In a practical use case, the approach outperforms classical MES solutions and combinations of MES with genetic algorithms in terms of makespan reduction.
@@
This paper by Carugno et al. [27] also investigates the application of QA to the JSP. The study provides a detailed analysis of applying QA to JSP, covering problem formulation, quantum annealer fine-tuning, and comparison with classical solvers. The research addresses often-overlooked aspects, such as computational costs, qubit requirements, and mitigating chain breaks. Additionally, advanced tools like reverse annealing are explored. Results reveal challenges at various stages, highlighting important research questions and areas for improvement in QA technology. The study contributes to a balanced discussion on QA’s potential and limitations in solving JSP, emphasising open research directions for further exploration.
Windmann et al. [162] introduce a QAOA based approach to job scheduling for automated storage and retrieval systems. It formulates the problem as an asymmetric QUBO and uses QAOA for solutions. Evaluations show promising results in minimising transport costs and improving scalability.
@@
A factory scheduling problem was presented by Yarkoni et al. [167], derived from a real-world optimisation problem involving re-ordering colour sequences in a paint shop during the manufacturing of cars. The task is to assign colours to sequences of cars such that all customer orders are fulfilled while simultaneously minimising the number of colour switches within the sequence. In this work, the authors investigated problem instances ranging from 10–3000 cars in a single sequence. The multi-car paint shop problem described in the paper has a simple Ising model representation, and instances derived directly from real-world data were solved using D-Wave’s QPUs, hybrid algorithms, and classical algorithms. The results showed that while the quantum hardware and hybrid algorithms were able to provide adequate solutions to the problem for smaller sizes, the performance of these algorithms approached that of a simple greedy algorithm in the large size limit. However, the analysis also concluded that QA is currently approaching the limit of industrially-relevant problem scales. Partly the same authors look at the same problem but then using the QAOA approach in [148]. This research combines numerical simulations with experimental data obtained from a trapped ion quantum computer, presenting a comparison between the performance of QAOA and classical heuristics, particularly in the infinite-size limit for noiseless quantum computation. The experimental findings show the quick degradation of the quantum algorithm’s efficiency as the problem size increases.
@@
Also Huang et al. [66] look at the paint shop problem. While previous research primarily targeted minimising color changeovers between vehicles, this study introduces a novel problem formulation considering not only color transitions but also the impact on product quality. By incorporating a machine learning model, which is not explained in the paper, to predict quality defects and leveraging quantum computing, the paper presents an approach for solving the generalised paint shop sequencing problem. The paper concludes with a case study showcasing the potential of QAOA in solving a small-scale sequencing problem, paving the way for future exploration in real-world, industrial-sized scenarios.
@@
The next problem coming from the aircraft industry, is the Aircraft Recovery Problem (ARP), a crucial aspect of airline disruption management. Given the financial pressures on airlines, aggravated by the recent pandemic, a swift and efficient solution to recover from disruptions is essential to minimise financial losses. Mori et al. [101] develop a Two-Stage algorithm to model ARP as a QUBO problem. Two experiments are conducted to evaluate the model’s performance using QC in comparison to classical solvers. The first experiment compares the proposed approach with a previous one [28] based on evolutionary computation algorithms, showing similar execution times and results. The second experiment compares a classical solver with a hybrid solver using the proposed model, demonstrating identical solutions but with faster execution times for the hybrid solver.
The thesis by Grange [57] focuses on designing and applying quantum algorithms to solve complex railway optimization problems. It explores both exact and heuristic quantum algorithms, including Variational Quantum Algorithms and Quantum Approximate Optimization Algorithms, to improve solution quality and computation time for combinatorial optimization challenges faced by railroad companies.
@@
A more generic problem that also has applications in the aircraft industry is the Quadratic Assignment Problem (QAP). Mohammadbagherpoor [98] explores the QAP in the application of airline gate-scheduling . Here, the QAP involves optimising the allocation of facilities (airline gates) to locations (flights) with known assignment costs. The paper introduces the concept of using VQE and space-efficient graph colouring, to enhance quantum computing algorithms for solving QAPs. The study tests these enhanced quantum computing algorithms on an 8-airline gate and 24-flight test case using both the IBM quantum computing simulator and a 27-qubit superconducting Transmon IBM quantum computing hardware platform. The results suggest that smaller circuit size and efficient mapping to hardware topology enable the application of variational quantum algorithms to larger problem sets, a critical advancement in solving optimisation problems.
@@
The work by Scherer et al. [135] discusses the application of Grover’s algorithm, to solve the “on-call spacecraft operator scheduling” problem at the German Space Operating Center (GSOC). The problem involves scheduling operators for spacecraft missions, considering various constraints and optimisation goals. The scheduling must allocate approximately 50 operators to 20 positions over 180 days, while satisfying constraints such as operator availability, working preferences, and limiting consecutive working days. Evaluation is done using Qiskit’s QASM simulator with up to 32 qubits. The success rates are evaluated for different problem sizes and configurations. The paper successfully demonstrates solving the scheduling problem using Grover’s algorithm on a quantum simulator. Future work includes testing on real quantum devices, improving qubit stability, and exploring quantum-classical hybrid approaches. Emphasis on minimising the number of qubits and optimising the algorithm’s width and depth is highlighted.
@@
Adelomou et al. [2] address personnel scheduling. They look at the challenges faced by social workers in generating optimal visiting schedules for patients. The problem involves a combination of routing and planning issues. The article suggest the use of the Parameterised Quantum Circuits (PQC) and the VQE algorithm in solving the problem. It explores the potential of PQC as a basis for Machine Learning (ML) and delves into the formulation of the social worker problem for quantum computing. The proposed methodology aims to offer an efficient and intelligent system, capable of learning by analogy and avoiding the need for recalculating from scratch. Experimental results suggest the effectiveness of the Quantum Cases-Based Reasoning (QCBR) approach, demonstrating potential benefits for optimising worker schedules. The article concludes by emphasising the broader applicability of the proposed formulation to solve various planning, scheduling, and routing problems. It envisions the creation of hybrid and intelligent systems that leverage quantum resources efficiently, reducing costs and laying the foundation for generalised problem-solving in the field.
Last, Krol et al. [79] show the design and implementation of a quantum algorithm for industrial shift scheduling (QISS), which uses Grover’s adaptive search to tackle a common and important class of valuable, real-world combinatorial optimization problems. They give an explicit circuit construction of the Grover’s oracle, incorporating the multiple constraints present in the problem, and detail the corresponding logical-level resource requirements. Further, they simulate the application of QISS to specific small-scale problem instances to corroborate the performance of the algorithm, and provide an open-source repository.